OpenStudy (anonymous):

Solve for y (4/7y)+(2/7)=6

OpenStudy (anonymous):

need common denominator: \[4/7y +2y/7y=(4+2y)/7y=6\] 4+2y = 6*7y can you finish it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

4+2y = 6*7y 40y=4 y=4/40 y=1/10

OpenStudy (anonymous):

(4/7y)+(2/7)=6 \[4/7y = 40/7\] \[4/y=40\] y=1/10